# Prospects for sub-EW ALP searches via $\gamma+b\bar{b}$ signatures at the LHC using jet substructure techniques

### v1

@@ 1 Introduction
The excellent performance of the LHC in terms of delivered luminosity has allowed the ATLAS and CMS experiments to set stringent limits on new particle masses well beyond the electroweak (EW) scale, thus worsening the naturalness problem. However, it is still possible that the new physics is hidden at lower energies, being weakly coupled to the known Standard Model (SM) particles, such that any sign of its production are swamped by the SM background. Current resonance searches at the LHC leave an area of opportunity roughly below the EW scale, which could potentially be accessed via the $b\bar{b}$ final state. Light scalars and pseudoscalars are particularly promising candidates, with possibly large branching ratios to $b\bar{b}$ if their masses are above 10 GeV, as further discussed in Section 2. The resonances may be identified with well-motivated axion-like particles (ALPs) Jaeckel:2010ni ; DiLuzio:2020wdo ; Brivio:2017ije ; Bauer:2017ris ; ParticleDataGroup:2024cfk . They could be mediators between the SM and dark matter (DM) XiaoyongChu_2012 ; Boehm:2014hva ; Arina:2014yna ; Hambye:2019dwd ; Bharucha:2022lty . Such spin-zero particles with masses below the EW scale could further arise in models such as the next-to-minimal supersymmetric SM Ellwanger:2009dp , or composite Higgs models with additional Nambu-Goldstone bosons Gripaios:2009pe ; Frigerio:2012uc ; Cacciapaglia:2019bqz ; BuarqueFranzosi:2021kky ; Elander:2020nyd .
@@
In this work, we present a sensitivity study for the search of pseudoscalar resonances decaying to $b\bar{b}$ pairs in proton-proton collisions at a center-of-mass energy of $\sqrt{s}=14$ TeV and an integrated luminosity of 3000 fb-1, achievable at the end of the High Luminosity LHC (HL-LHC) data taking, using a parametrised ATLAS detector simulation. We perform our analysis focusing on the sensitivity for $b\bar{b}$ resonances with mass values ranging from 10 GeV to 100 GeV. In this mass range, we adopt a strategy based on triggering on a high-$p_{\rm T}$ photon and we look for a recoiling large-radius jet that contains collimated $b\bar{b}$ pairs. As discussed before, this is an optimal strategy, given the lower achievable photon threshold at trigger level with respect to choosing ISR jets, it can probe efficiently masses $\gtrsim 10$ GeV. Moreover, the $b\bar{b}$ final state is the dominant decay channel for spin-0 particles with Yukawa-like couplings to fermions and masses below the top-quark mass. To be noticed that the analysis proposed is an extension of the current landscape of light $b\bar{b}$ resonance searches.
@@ 2 Light pseudoscalar: model and motivations
$\mathcal{L}=$ $\;{\cal L}_{\rm SM}+{\mathcal{L}}_{\rm ALP}+\bar{\chi}(i\partial\!\!\!/-m_{\chi})\chi+i\frac{C_{\chi}}{f_{a}}\,m_{\chi}\,a\bar{\chi}\gamma_{5}\chi\,,$ (3)
where the DM particle $\chi$ of mass $m_{\chi}$ couples to the ALP via the coupling $C_{\chi}$. Depending on the couplings of the ALP to DM and to the SM, the production mechanism changes. In freeze-out, one has large enough ALP couplings such that equilibrium is obtained between the hidden sector particles and SM, whereas in freeze-in, the couplings are small, such that DM-SM equilibrium is never achieved. In both cases, the relic density may depend on the ALP to DM thermal cross-section or the SM to DM thermal cross-section, depending on which of these are dominant. Considering the $m_{a}$ $g_{\rm aff}$ plane, it was found in Ref. Bharucha:2022lty that a large range of parameter space, where the correct relic density via freeze-in can be obtained, remains allowed, particularly for values of $m_{a}$ above $\mathcal{O}(10)$ GeV. Note that for the values of $g_{\rm aff}$ close to the theoretical upper limit of $g_{\rm aff}\lesssim 0.01$ GeV-1 (which could be probed at colliders) dark matter production via fermion scattering is relevant. It was seen in Ref. Bharucha:2022lty that in this region there is a large hierarchy between $g_{\rm aff}$ and $g_{a\chi\chi}$, which would need to be explained via further model building.
The sensitivity study performed in this work is presented in terms of the above-discussed ALP model. However, pseudoscalars with masses below the electroweak scale and couplings to fermions are also motivated by theories addressing the stability of the electroweak scale, such as Higgs compositeness. In fact, a composite Higgs boson requires the introduction of a new strongly-coupled sector close to the TeV scale, whose spectrum contains several composite resonances, including SM singlets that may be lighter than the Higgs itself. In this context, a particularly well-motivated target for low mass resonance searches is a pseudo Nambu-Goldstone boson $a$, analogous to a QCD pion, whose small mass arises radiatively, via shift-symmetry breaking couplings. Indeed, the couplings of the $a$ to $b$-quarks induce $m_{a}\gtrsim 10$ GeV Gripaios:2009pe ; Frigerio:2012uc at one loop, as desired. Note that here we are referring to additional ALP couplings, which break the Nambu-Goldstone shift symmetry, thus inducing a non-trivial potential. Such a composite pseudoscalar has been searched for at colliders, but mostly using large-$p_{\rm T}$ signatures which require $m_{a}$ above the electroweak scale Cacciapaglia:2019bqz . An interesting study to probe $m_{a}$ well below the electroweak scale at LHCb can be found in Ref. BuarqueFranzosi:2021kky , which exploits decay channels other than $b\bar{b}$.
@@ 3 Collider Analysis
The dominant background arises from multijet production in association with a photon. In Fig. 1 (b), we show one of the Feynman diagrams for the $\gamma+$multijet process. Subdominant resonant backgrounds originate from $W+\gamma$ and $Z+\gamma$ processes (see Fig. 1 (c)). In the following subsections, we discuss the generation of signal and background processes, the physics object used in the analysis, kinematic observables, the event selection criteria and finally the details of the $b$-tagging simulation. The section ends with the estimation of the signal and background yields.
@@ 3.1 Event simulation
Signal and SM background processes are simulated using Monte Carlo (MC) event generators. The $\gamma+$ALP signal process $pp\to a(\to b\bar{b})\gamma$ is generated using MadGraph5_aMC@NLO Alwall:2014hca at leading order (LO). The $\gamma$+multijet background process, which include both non-resonant and resonant $W/Z+\gamma$ processes, is simulated using Pythia8 Sjostrand:2014zea . Both signal and background are generated in the 5F scheme where proton contains gluon, light- and $b$-quarks. The center-of-mass energy considered for the proton-proton collisions is $\sqrt{s}=14$ TeV. To simulate the showering and hadronisation effects, the generated events are interfaced with Pythia8, with the A14 tune ATL-PHYS-PUB-2014-021 , and the parton distribution function (PDF) set NNPDF2.3NLO NNPDF:2014otw is used. Finally, the detector simulation is performed with Delphes-3.5.0 deFavereau:2013fsa using the default HL-LHC ATLAS analysis card. We apply specific requirements for the generation of the signal and background events to improve the statistics in the phase space relevant to the proposed analysis. The inclusive cross-sections and generation-level cut efficiencies for the $pp\to a(\to b\bar{b})\gamma$ signal process is shown in Table 1.
@@
$m_{a}$ $\rm[GeV]$ $\sigma$ ($g_{\mathrm{aff}}=0.1$ GeV-1) $\rm[pb]$ $\epsilon$ (×$\times 10^{-5}$)
@@ 3.2 Object definitions and event selection
At selection level, a large radius jet for full containment of the hadronisation products, coming from the final state $b$-quarks, is required. The reconstruction of jets is performed within the FastJet framework Cacciari:2011ma , utilizing the AK1.0 anti-$k_{\textrm{T}}$ clustering algorithm, using the particle-flow objects in Delphes as input. The photon, used to trigger on, is required to have $p_{\mathrm{T},\gamma}>160$ GeV and $|\eta_{\gamma}|<2.1$ ATLAS:2018hbc ; CMS:2019xai . Further, a separation between the photon and AK1.0 jet is required to be $\Delta R>2.2$  [^2] The angular distance in the pseudorapidity ($\eta$) and azimuthal angle ($\phi$) plane is measured in units of $\Delta R=\sqrt{(\Delta\eta)^{2}+(\Delta\phi)^{2}}$. Here, $\eta$ is defined in terms of polar angle ($\theta$) as $\eta=-{\rm ln~{}tan}(\theta/2)$. CMS:2019xai . We apply the soft drop mass algorithm Larkoski:2014wba to groom the AK1.0 jet, using the parameters $z_{cut}$ = 0.1 and $\beta$ = 1 ATL-PHYS-PUB-2021-028 where, $z_{cut}$ denotes the energy threshold required to satisfy the soft drop condition during the declustering procedure. The soft drop condition is defined as $\frac{\text{min}(p_{\mathrm{T},1},p_{\mathrm{T},2})}{p_{\mathrm{T},1}+p_{\mathrm{T},2}}>z_{cut}~{}(\frac{\Delta R_{12}}{R_{0}})^{\beta}$. Here, $p_{\mathrm{T},i}$ are jet constituents’ transverse momentum, $\Delta R_{12}$ is the angular separation between those constituents and $R_{0}$ is the jet radius parameter. The angular exponent $\beta\neq 0$ imposes an additional angular requirement on the soft drop condition. Specifically, the soft drop removes soft wide angle radiation from a jet, while keeping a fraction of the soft collinear radiation as long as $\beta>0$. Although the $\beta=0$ choice is widely used CMS:2019xai , this choice significantly reduces the signal efficiency the for lowest ALP mass ($m_{a}$ = 12 GeV) considered in this analysis while $\beta=1$ is found to be optimal across the entire mass range. Fig. 2 represents the distribution of $\mathrm{m}_{\mathrm{SD}}$, the invariant mass of the soft dropped AK1.0 jet. As shown, by removing the soft radiation, the grooming technique soften the $\mathrm{m}_{\mathrm{SD}}$ spectrum of multijet events, while it leaves the signal distribution almost unaltered.
@@
Here, $z_{i}$ is the energy fraction of $i$th jet constituent, $z_{i}\equiv\frac{p_{\mathrm{T},i}}{p_{\mathrm{T},\mathrm{jet}}}$ and $\Delta R_{ij}$ corresponds to the separation between the $i$th and $j$th jet constituents in the $\eta-\phi$ plane. In general, ${}_{v}e_{n}$ depends on the $n$ energy fractions of the $N$ constituent particles inside the jet ($n\leq N$), relative to the jet $p_{\rm T}$ and the minimal product of $v$ number of pairwise angles between the $n$ selected particles. In a scenario where there is a perfect two-prong structure inside the jet, $(_{1}e_{2})^{2}\gg~{}_{2}e_{3}$ Moult:2016cvt , therefore the smaller value of $\mathrm{N}_{2}$ in signal events can be used as discriminant against multijet background. At analysis level, in order to reduce the variable’s dependency on the jet mass and transverse momentum, the $\mathrm{N}_{2}$ variable is estimated on the groomed AK1.0 jet.
Figure 3: The $10\%$ quantile of the $\mathrm{N}_{2}$ distribution in simulated multijet events, $\mathrm{N}_{2}^{10\%}$, used to define the $\mathrm{N}_{2}^{\mathrm{DDT}}$ variable. The distribution is shown in the $\rho-p_{\rm T}$ plane. The white dashed lines correspond to constant values of the jet mass, $\mathrm{m}_{\mathrm{SD}}$. We select the region $-7.5<\rho<-2.0$ for our analysis, as shown with vertical gray lines.
The dimensionless scaling variable $\rho$ defined as $\rho=\mathrm{log}(\mathrm{m}_{\mathrm{SD}}^{2}/p_{\rm T}^{2})$ Dolen:2016kst is designed to reject events having non-perturbative effects in the QCD jet mass spectrum. The selected $\rho$ region in the analysis is $-7.5\leq\rho\leq-2.0$, which is efficient for the signal process, and it corresponds to our region of interest, with $\mathrm{m}_{\mathrm{SD}}$ from 5 GeV to 100 GeV, for AK1.0 jet with $p_{\rm T}$ = 200 GeV. $\mathrm{N}_{2}^{10\%}$, shown in Fig. 3, is evaluated in multiple bins of the $\rho-p_{\rm T}$ plane. Smoothing, using the distance-weighted $k$-nearest-neighbor (kNN) rule 5408784 , is applied to $\mathrm{N}_{2}^{10\%}$, to efficiently populate all the phase space. The signal efficiency of the $\mathrm{N}_{2}^{\mathrm{DDT}}<0$ requirement, varies roughly between $12\%$ and $26\%$, increasing the signal over background ratio, since by construction only 10% of the background are retained, with the exception of $m_{a}=12$ GeV for which the efficiency drops to $\sim 1\%$. For the latter, given the high $p_{\rm T}$ of the AK1.0 jet, the $b$-quarks from the ALP decay are so collimated that the two-prong structure is lost, resulting in a large loss in efficiency after the $\mathrm{N}_{2}^{\mathrm{DDT}}$ requirement. We will show later that this can be partly recovered if a lower threshold for the trigger photon can be achieved by the LHC experiments.
@@
$p_{\mathrm{T},\gamma}$ $>160$ GeV
$|\eta_{\gamma}|$ $<2.1$
$p_{\mathrm{T,AK1.0}}$ $>200$ GeV
[…]
$\rho$ $[-7.5,-2.0]$
$\mathrm{N}_{2}^{\mathrm{DDT}}$ $<0$
[…]

### v3

# Prospects for sub-EW-scale ALP searches via $\gamma+b\bar{b}$ signatures at the LHC using jet substructure techniques

@@ 1 Introduction
The excellent performance of the LHC in terms of delivered luminosity has allowed the ATLAS and CMS experiments to set stringent limits on new particle masses well beyond the electroweak (EW) scale, thus worsening the naturalness problem. However, it is still possible that the new physics is hidden at lower energies, being weakly coupled to the known Standard Model (SM) particles, such that any sign of its production are swamped by the SM background. Current resonance searches at the LHC leave an area of opportunity roughly below the EW scale, which could potentially be accessed via the $b\bar{b}$ final state. Light scalars and pseudoscalars are particularly promising candidates, with possibly large branching ratios to $b\bar{b}$ if their masses are above 10 GeV, as further discussed in Section 2. The resonances may be identified with well-motivated axion-like particles (ALPs) Jaeckel:2010ni ; DiLuzio:2020wdo ; Brivio:2017ije ; Bauer:2017ris ; ParticleDataGroup:2024cfk . They could be mediators between the SM and dark matter (DM) XiaoyongChu_2012 ; Boehm:2014hva ; Arina:2014yna ; Hambye:2019dwd ; Bharucha:2022lty . Such spin-zero particles with masses below the EW scale could further arise in supersymmetric models, such as the next-to-minimal supersymmetric SM Ellwanger:2009dp or models with an approximate $R$-symmetry Bellazzini:2017neg , or composite Higgs models with additional Nambu-Goldstone bosons Gripaios:2009pe ; Frigerio:2012uc ; Cacciapaglia:2019bqz ; BuarqueFranzosi:2021kky ; Elander:2020nyd .
@@
In 2030 the High Luminosity (HL-LHC) phase of the LHC will begin ATL-PHYS-PUB-2019-005 ; Collaboration:2650976 , which will allow the ATLAS and CMS experiments to collect an integrated luminosity of 3000 fb-1 of proton-proton collisions at a center-of-mass energy of $\sqrt{s}=14$ TeV. Right before HL-HLC starts, both detectors will undergo several upgrades that will result in improved reconstruction and pile-up suppression capabilities. This work aims at optimizing the analysis strategy in order to maximize the experimental sensitivity to detect pseudoscalar resonances decaying to $b\bar{b}$ pairs at the end of HL-HLC. We perform our analysis focusing on the sensitivity for resonances with mass values ranging from 10 GeV to 100 GeV. In this mass range, we adopt a strategy based on triggering on a high-$p_{\rm T}$ photon and we look for a recoiling large-radius jet that contains collimated $b\bar{b}$ pairs. As discussed before, this is an optimal strategy, given the lower achievable photon threshold at trigger level with respect to choosing ISR jets, it can probe efficiently masses $\gtrsim 10$ GeV. Moreover, the $b\bar{b}$ final state is the dominant decay channel for spin-0 particles with Yukawa-like couplings to fermions and masses below the top-quark mass. To be noticed that the analysis proposed is an extension of the current landscape of light $b\bar{b}$ resonance searches.
@@ 2 Light pseudoscalar: model and motivations
$\mathcal{L}=$ $\;{\cal L}_{\rm SM}+{\mathcal{L}}_{\rm ALP}+\bar{\chi}(i\partial\!\!\!/\penalty 50-m_{\chi})\chi+i\frac{C_{\chi}}{f_{a}}\,m_{\chi}\,a\bar{\chi}\gamma_{5}\chi\,$, (3)
where the DM particle $\chi$ of mass $m_{\chi}$ couples to the ALP via the coupling $C_{\chi}$. Depending on the couplings of the ALP to DM and to the SM, the production mechanism changes. In freeze-out, one has large enough ALP couplings such that equilibrium is obtained between the hidden sector particles and SM, whereas in freeze-in, the couplings are small, such that DM-SM equilibrium is never achieved. In both cases, the relic density may depend on the ALP to DM thermal cross-section or the SM to DM thermal cross-section, depending on which of these are dominant. Considering the $m_{a}$ - $g_{\rm aff}$ plane, it was found in Ref. Bharucha:2022lty that a large range of parameter space, where the correct relic density via freeze-in can be obtained, remains allowed, particularly for values of $m_{a}$ above $\mathcal{O}(10)$ GeV. Note that for the values of $g_{\rm aff}$ close to the theoretical upper limit of $g_{\rm aff}\lesssim 0.01$ GeV-1 (which could be probed at colliders) dark matter production via fermion scattering is relevant. It was seen in Ref. Bharucha:2022lty that in this region there is a large hierarchy between $g_{\rm aff}$ and $g_{a\chi\chi}$, which would need to be explained via further model building.
The sensitivity study performed in this work is presented in terms of the above-discussed ALP model. However, pseudoscalars with masses below the electroweak scale and couplings to fermions are also motivated by theories addressing the stability of the electroweak scale, such as Higgs compositeness. In fact, a composite Higgs boson requires the introduction of a new strongly-coupled sector close to the TeV scale, whose spectrum contains several composite resonances, including SM singlets that may be lighter than the Higgs itself. In this context, a particularly well-motivated target for low mass resonance searches is a pseudo Nambu-Goldstone boson $a$, analogous to a QCD pion, whose small mass arises radiatively, via shift-symmetry breaking couplings. Indeed, the couplings of the $a$ to $b$-quarks induce $m_{a}\gtrsim 10$ GeV Gripaios:2009pe ; Frigerio:2012uc at one loop, as desired. Note that here we are referring to additional ALP couplings, which break the Nambu-Goldstone shift symmetry, thus inducing a non-trivial potential. Such a composite pseudoscalar has been searched for at colliders, but mostly using large-$p_{\rm T}$ signatures which require $m_{a}$ above the electroweak scale Cacciapaglia:2019bqz . Interesting studies to probe $m_{a}$ well below the electroweak scale at LHCb can be found in Refs. CidVidal:2018blh ; BuarqueFranzosi:2021kky , which exploit decay channels other than $b\bar{b}$.
@@ 3 Collider Analysis
The dominant background arises from multijet production in association with a photon. In Fig. 1 (b), we show one of the Feynman diagrams for the $\gamma+$multijet process. Subdominant resonant backgrounds originate from $W+\gamma$ and $Z+\gamma$ processes (see Fig. 1 (c)). Since modeling uncertainties could have a large impact on the results, simulated background events are only used here to validate the strategy of the data-driven background estimation. Despite a plethora of next-to-leading order (NLO) generators Catani:2002ny ; Sherpa:2024mfk being available and currently used by both ATLAS and CMS collaborations to compare with data in specific phase-space regions ATLAS:2016fta ; ATLAS:2023yrt , the LHC experiments still rely on leading order (LO) generators to provide large samples of $\gamma$ + jets events to validate data-driven techniques CMS:2019xai ; CMS:2017dcz . In the following subsections, we discuss the generation of signal and background processes, the physics object used in the analysis, kinematic observables, the event selection criteria, and finally the details of the $b$-tagging simulation. The section ends with the estimation of the signal and background yields.
@@ 3.1 Event simulation
Signal and SM background processes are simulated using Monte Carlo (MC) event generators. The $\gamma+$ALP signal process $pp\to a(\to b\bar{b})\gamma$ is generated using MadGraph5_aMC@NLO Alwall:2014hca at LO. The $\gamma$+multijet background process, which include both non-resonant and resonant $W/Z+\gamma$ processes, is simulated using LO generator Pythia8 Sjostrand:2014zea . Both signal and background are generated in the 5F scheme where proton contains gluon, light- and $b$-quarks. The center-of-mass energy considered for the proton-proton collisions is $\sqrt{s}=14$ TeV. To simulate the showering and hadronisation effects, the generated events are interfaced with Pythia8, with the A14 tune ATL-PHYS-PUB-2014-021 , and the parton distribution function (PDF) set NNPDF2.3NLO NNPDF:2014otw is used. Finally, the detector simulation is performed with Delphes-3.5.0 deFavereau:2013fsa using the default HL-LHC ATLAS analysis card. We apply specific requirements for the generation of the signal and background events to improve the statistics in the phase space relevant to the proposed analysis. The inclusive cross-sections and generation-level cut efficiencies for the $pp\to a(\to b\bar{b})\gamma$ signal process is shown in Table 1.
@@
$m_{a}$ $\rm[GeV]$
$\sigma$ ($g_{\mathrm{aff}}=0.1$ GeV-1) $\rm[pb]$
$\epsilon$ ($\times 10^{-5}$)
@@ 3.2 Object definitions and event selection
At selection level, a large radius jet for full containment of the hadronisation products, coming from the final state $b$-quarks, is required. The reconstruction of jets is performed within the FastJet framework Cacciari:2011ma , utilizing the AK1.0 anti-$k_{\textrm{T}}$ clustering algorithm, using the particle-flow objects in Delphes as input. The photon, used to trigger on, is required to have $p_{\mathrm{T},\gamma}>160$ GeV and $|\eta_{\gamma}|<2.1$ ATLAS:2018hbc ; CMS:2019xai . Further, a separation between the photon and AK1.0 jet is required to be $\Delta R>2.2$  [^2] The angular distance in the pseudorapidity ($\eta$) and azimuthal angle ($\phi$) plane is measured in units of $\Delta R=\sqrt{(\Delta\eta)^{2}+(\Delta\phi)^{2}}$. Here, $\eta$ is defined in terms of polar angle ($\theta$) as $\eta=-{\rm ln~tan}(\theta/2)$. CMS:2019xai . We apply the soft drop mass algorithm Larkoski:2014wba to groom the AK1.0 jet, using the parameters $z_{cut}$ = 0.1 and $\beta$ = 1 ATL-PHYS-PUB-2021-028 where, $z_{cut}$ denotes the energy threshold required to satisfy the soft drop condition during the declustering procedure. The soft drop condition is defined as $\frac{\text{min}(p_{\mathrm{T},1},p_{\mathrm{T},2})}{p_{\mathrm{T},1}+p_{\mathrm{T},2}}>z_{cut}~(\frac{\Delta R_{12}}{R_{0}})^{\beta}$. Here, $p_{\mathrm{T},i}$ are jet constituents’ transverse momentum, $\Delta R_{12}$ is the angular separation between those constituents and $R_{0}$ is the jet radius parameter. The angular exponent $\beta\neq 0$ imposes an additional angular requirement on the soft drop condition. Specifically, the soft drop removes soft wide angle radiation from a jet, while keeping a fraction of the soft collinear radiation as long as $\beta>0$. Although the $\beta=0$ choice is widely used CMS:2019xai , this choice significantly reduces the signal efficiency the for lowest ALP mass ($m_{a}$ = 12 GeV) considered in this analysis while $\beta=1$ is found to be optimal across the entire mass range. Fig. 2 represents the distribution of $\mathrm{m}_{\mathrm{SD}}$, the invariant mass of the soft dropped AK1.0 jet. As shown, by removing the soft radiation, the grooming technique soften the $\mathrm{m}_{\mathrm{SD}}$ spectrum of multijet events, while it leaves the signal distribution almost unaltered. This, in turn, presents the advantage of allowing the background shape to be fitted using an exponentially falling distribution, which can be constrained by data on both sides of the resonant mass values (side-band method ATL-PHYS-PUB-2020-028 ). This is shown in Fig. 2 (a) for the case $\beta=1.0$, where the $\mathrm{m}_{\mathrm{SD}}$ distribution for multijet events presents an exponentially falling behaviour starting from values as low as 10 GeV. In this particular case, a deviation can be seen starting from $\mathrm{m}_{\mathrm{SD}}=80$ GeV, consistent with resonant W- and Z-boson production.
@@
Here, $z_{i}$ is the energy fraction of $i$th jet constituent, $z_{i}\equiv\frac{p_{\mathrm{T},i}}{p_{\mathrm{T},\mathrm{jet}}}$ and $\Delta R_{ij}$ correspond to the separation between the $i$th and $j$th jet constituents in the $\eta-\phi$ plane. In general, $_{v}e_{n}$ depends on the $n$ energy fractions of the $N$ constituent particles inside the jet ($n\leq N$), relative to the jet $p_{\rm T}$ and the minimal product of $v$ number of pairwise angles between the $n$ selected particles. In a scenario where there is a perfect two-prong structure inside the jet, $(_{1}e_{2})^{2}\gg~_{2}e_{3}$ Moult:2016cvt , therefore the smaller value of $\mathrm{N}_{2}$ in signal events can be used as discriminant against multijet background. At analysis level, in order to reduce the variable’s dependency on the jet mass and transverse momentum, the $\mathrm{N}_{2}$ variable is estimated on the groomed AK1.0 jet.
Figure 3: The $10\%$ quantile of the $\mathrm{N}_{2}$ distribution in simulated multijet events, $\mathrm{N}_{2}^{10\%}$, used to define the $\mathrm{N}_{2}^{\mathrm{DDT}}$ variable. The distribution is shown in the $\rho-p_{\rm T}$ plane. The white dashed lines correspond to constant values of the jet mass, $\mathrm{m}_{\mathrm{SD}}$. We select the region $7.5<\rho<-2.0$ for our analysis, as shown with vertical gray lines.
The dimensionless scaling variable $\rho$ defined as $\rho=\mathrm{log}(\mathrm{m}_{\mathrm{SD}}^{2}/p_{\rm T}^{2})$ Dolen:2016kst is designed to reject events having non-perturbative effects in the QCD jet mass spectrum. The selected $\rho$ region in the analysis is $7.5\leq\rho\leq-2.0$, which is efficient for the signal process, and it corresponds to our region of interest, with $\mathrm{m}_{\mathrm{SD}}$ from 5 GeV to 100 GeV, for AK1.0 jet with $p_{\rm T}$ = 200 GeV. $\mathrm{N}_{2}^{10\%}$, shown in Fig. 3, is evaluated in multiple bins of the $\rho-p_{\rm T}$ plane. Smoothing, using the distance-weighted $k$-nearest-neighbor (kNN) rule 5408784 , is applied to $\mathrm{N}_{2}^{10\%}$, to efficiently populate all the phase space. The signal efficiency of the $\mathrm{N}_{2}^{\mathrm{DDT}}<0$ requirement, varies roughly between $12\%$ and $26\%$, increasing the signal over background ratio, since by construction only 10% of the background are retained, with the exception of $m_{a}=12$ GeV for which the efficiency drops to $\sim 1\%$. For the latter, given the high $p_{\rm T}$ of the AK1.0 jet, the $b$-quarks from the ALP decay are so collimated that the two-prong structure is lost, resulting in a large loss in efficiency after the $\mathrm{N}_{2}^{\mathrm{DDT}}$ requirement. We will show later that this can be partly recovered if a lower threshold for the trigger photon can be achieved by the LHC experiments.
@@
$p_{\mathrm{T},\gamma}$ >160>160 GeV
$|\eta_{\gamma}|$ <2.1<2.1
$p_{\mathrm{T,AK1.0}}$ >200>200 GeV
$\Delta R(\gamma,~\text{AK1.0})$ >2.2>2.2
$\rho$ [7.5,2.0][-7.5,-2.0]
$\mathrm{N}_{2}^{\mathrm{DDT}}$ <0<0
Table 2: Summary of the selection requirements. The transverse momentum of the photon and AK1.0 jet is denoted as $p_{\mathrm{T},\gamma}$ and $p_{\mathrm{T,AK1.0}}$, respectively, and distance between them is referred as $\Delta R(\gamma,~\text{AK1.0})$. $\eta_{\gamma}$ is the pseudorapidity of the photon. $\rho$ is a dimensionless scaling variable, $\rho=\mathrm{log}(\mathrm{m}_{\mathrm{SD}}^{2}/p_{\rm T}^{2})$. The DDT observable is represented as $\mathrm{N}_{2}^{\mathrm{DDT}}$.